The length of the edge of the square is twice as long as the radius of the circle. Find the ratio of the area of the circle to the area of the square. Write the answer in simplified form.
Answer by mananth(16946)

circle radius = r
area = pir^2
square side = 2r
area = 4r^2
area of circle / area of square = pir^2/4r^2
= pi/4